OpenStudy (anonymous):

integral of (4x-2)^3

OpenStudy (anonymous):

u-sub, did you try that?

OpenStudy (anonymous):

oops my problem is actually integral of (4-2x)^3

OpenStudy (anonymous):

sorry :| but yes @sourwing but some how I'm getting stuck on this particular integral

OpenStudy (anonymous):

i am getting the answer -(1/8)(4-2x)^4 but my book has another answer

OpenStudy (atlas):

there are two ways you can do this - the lengthy way - break (4-2x)^3 using (a-b)^3 and apply integral over each part

OpenStudy (anonymous):

that's correct.

OpenStudy (atlas):

the short way - use substitution by assuming y = 4-2x

OpenStudy (anonymous):

what does the book have?

OpenStudy (anonymous):

the book has...

OpenStudy (anonymous):

can you simplify (-1/8) (4-2x)^4 ?

OpenStudy (anonymous):

(-1/8) (4-2x)^4 = (-1/8)(2)^4 (2-x)^4 = -2(2-x)^4

OpenStudy (anonymous):

your book simplified the integral. The two are still the same

OpenStudy (atlas):

@sourwing right: they just took the 2 as common from the part (4-2x) to write 2(2-x)

OpenStudy (anonymous):

ahhh i figured that was the case i just wasn't totally sure if i could pull out the 2!